Suppose your business has a 40% chance of making $40 and a 60% chance of making $100.

If the Utility function is the Square Root of Income, how much would you be willing to pay in insurance PREMIUMS in order to get the same level of Utility (happiness) with NO risk vs. a bunch of risk (variability) at your Expected Income?
Answer #1

SOLUTION:-

* Expected utility = 0.4 Utility(40) + 0.6 Utility(100)

* Expected utility = 0.4(40)0.5 + 0.6(100)0.5=3.16227766+6=9.16227766

* Thus in order to get same utility, incime required=83.9473

* Thus insurance amount=100-84.973=16.05

* Thus insurance premium=16.05

THANK YOU

If any quearies please leave your valuable comment on comment box......